Consider the following numerical example of the simple Kenesian model with no government spending,taxes or a foreign sector (all figures in R millions):
C =100 + 0,9Y
I = 50
Answer the following questions.
1. What is the value of the marginal propensity to consume(MPC) in this model?(2 marks)
2. Use a graph to illustrate the equilibrium level of output (5 marks)
3. Calculate the equilibrium level of output. (5 marks)
1
Expert's answer
2019-08-08T10:59:54-0400

1. The value of the marginal propensity to consume is: MPC = 0.9.

2. The equilibrium level of output is shown as an intersection point of AD and AS(Y) curves.

3. The equilibrium level of output is:

Y = C + I = 100 + 0.9Y + 50,

0.1Y = 150,

Y = 1500.